# Generative Modeling with Manifold Percolation: Metrics, Dynamics, and Topological Supervision

### v1

@@ Abstract
Generative modeling is typically framed as learning mapping rules, but from an observer’s perspective without access to these rules, the task manifests as disentangling the geometric support from the probability distribution. We propose that Continuum Percolation is uniquely suited for this support analysis, as the sampling process effectively projects high-dimensional density estimation onto a geometric counting problem on the support. In this work, we establish a rigorous isomorphism between the topological phase transitions of Random Geometric Graphs and the underlying data manifold in high-dimensional space. By analyzing the relationship between our proposed Percolation Shift metric and FID, we demonstrate that our metric captures structural pathologies (such as implicit mode collapse) where statistical metrics fail. Finally, we translate this topological phenomenon into a differentiable loss function to guide training. Experimental results confirm that this approach not only prevents manifold shrinkage but drives the model toward a state of ”Hyper-Generalization,” achieving good fidelity and verified topological expansion.
@@ 1 Theoretical Framework: Manifold Percolation
In this section, we establish the mathematical foundation of our framework. We bridge the gap between generative modeling and statistical physics by formalizing the ”Observer’s View” through the lens of Continuum Percolation Theory.
@@ 1.3 Scaling Laws and Intrinsic Dimension
Solving for $\varepsilon_{c}$ at $\langle k\rangle=\lambda_{c}$ yields $\varepsilon_{c}\propto(\int p^{2}(x)dx)^{-1/d}N^{-1/d}$. This confirms that the scaling law is robust to density variations, with the ”effective volume” determined by the Rényi entropy. ∎
@@ 1.5 Metric Spaces: From Pixel to Semantic Percolation
[…]
This projection allows us to detect ”Semantic Collapse,” where a model might maintain pixel variance (noise) but fails to cover the semantic diversity of the true manifold.
@@ 2 Manifold Dynamics and Topological Diagnosis
$\Delta\varepsilon_{c}=\varepsilon_{c}(\text{Model})-\varepsilon_{c}(\text{Real})$ (12)
According to Theorem 1, a negative shift ($\Delta\varepsilon_{c}<0$) serves as a rigorous proxy for manifold shrinkage ($\mathrm{Vol}(\mathcal{M}_{\text{model}})<\mathrm{Vol}(\mathcal{M}_{\text{real}})$).
@@ 2.1 The Geometric Mechanism of Shrinkage
To validate our theoretical framework, we first conduct a controlled experiment on a synthetic 2D manifold. We simulate ”Implicit Mode Collapse” by reducing the intra-mode variance of a Gaussian mixture while preserving the mode centers. As predicted, the percolation phase transition curve shifts to the left, resulting in a negative $\Delta\varepsilon_{c}$. This confirms that our metric is sensitive to the effective volume of the support, effectively distinguishing between a dispersed (healthy) manifold and a contracted (collapsed) one.
@@
Figure 2: Controlled Experiment on a Toy Manifold. (a) Simulating Implicit Mode Collapse by reducing variance (red) relative to ground truth (gray). (b) The percolation curve shifts left ($\Delta\varepsilon_{c}<0$), validating Theorem 1.
2.2 The Anatomy of Implicit Mode Collapse
Moving to real-world data (CIFAR-10), we perform a comprehensive diagnostic across training epochs (Ep 0, 5, 60, 100, 300). By integrating feature space visualization, multi-view percolation analysis, and training dynamics, we uncover a critical ”illusion” in generative evaluation. As illustrated in Figure 3, our analysis reveals three interconnected phenomena:
  • Visual Contraction (Fig 3a): UMAP projections reveal a clear ”Expansion-then-Contraction” dynamic. While the model initially learns to span the space (Ep 60-100), the support physically contracts in the late training stage (Ep 300, Red), where samples cluster tightly, leaving visible voids compared to the reference topology.
  • The Pixel Space Illusion (Fig 3b, left): In raw pixel space, the percolation curves for trained models (Ep 60-300) overlap perfectly with Real Data. This suggests that the model successfully saturates the signal space, explaining why standard metrics (sensitive to pixel statistics) fail to detect issues.
  • The Semantic Gap (Fig 3b, right): In the semantic feature space, however, a persistent topological deficit remains. Despite training progress, the curve for Epoch 300 (Yellow) remains strictly to the left of the Real Data baseline ($\Delta=-1.52$). This indicates that the model hits a ”Topological Glass Ceiling”—it fails to naturally expand enough to cover the full semantic diversity of the true distribution.
(a) Visualizing Manifold Dynamics: UMAP projections show the progression from noise (Ep 0) to expansion (Ep 60-100), and finally to Manifold Shrinkage at Epoch 300 (Red), where the distribution becomes noticeably denser with visible voids.
(b) Metric Decoupling: Pixel Space (left sub-panel) creates an illusion of convergence. In contrast, Semantic Space (right sub-panel) reveals the truth: even the best model (Ep 300) suffers from a topological deficit ($\Delta=-1.52$).
(c) Dynamics: Standard metrics like FID (Red) saturate early, failing to capture the late-stage shrinkage. Our Percolation Shift (Blue) acts as a sensitive probe for this structural pathology.
Figure 3: The Anatomy of Implicit Mode Collapse. We present a unified diagnostic framework. (a) Visually, the manifold contracts in late training. (b) Topologically, this is explained by the inability of the model to bridge the ”Semantic Gap” in feature space. (c) Statistically, this phenomenon is invisible to FID but clearly detected by our Percolation Shift.
2.3 Quantitative Verification
The visual evidence is corroborated by quantitative metrics (Table 1). While FID remains relatively low at Epoch 300 (indicating high image quality), the Percolation Shift remains negative (1.12-1.12), confirming that the generated manifold is topologically smaller than the ground truth.
Table 1: Evidence of persistent manifold shrinkage in modern diffusion models. On CIFAR-10 unconditional (N=1000N{=}1000, semantic space), FID improves dramatically from Epoch 0 to 300, yet the generated manifold first over-expands (+11.49) and then consistently under-covers the true support ($\Delta\varepsilon_{c}\approx-1.2$) throughout late training. More negative $\Delta\varepsilon_{c}$ indicates more severe shrinkage.
@@ 2.3 Quantitative Verification
Model State FID $\downarrow$ Shift ($\Delta\varepsilon_{c}$) Topology Status
Real Data 0.00 0.00 Reference
Epoch 0 41.02 11.49 Noise Init
Epoch 5 1.93 -1.44 Early shrinkage
Epoch 60 0.80 -1.15 Shrinkage
Epoch 300 0.62 -1.12 moderate Shrinkage
Figure 4: Evolution of Semantic Percolation Shift ($\Delta\varepsilon_{c}$) over Training Epochs. Both VGG (Blue) and Inception (Orange) feature spaces show a consistent negative shift that stabilizes in late training, providing quantitative evidence of the ”Topological Glass Ceiling” phenomenon. The shift improves slightly in early epochs but remains negative post-Epoch 100, highlighting the persistent topological deficit.
2.4 Discussion: The Ricci Flow Analogy
Our findings suggest that ”Mode Collapse” in diffusion models is often implicit. The model does not necessarily drop discrete modes, but instead suffers from Manifold Shrinkage.
Remark 1 (Theoretical Insight).
This shrinkage can be physically interpreted through geometric flow. The denoising objective acts similarly to Ricci Flow ($\partial_{t}g_{ij}=-2R_{ij}$), which smooths manifold geometry to minimize curvature. Excessive smoothing during late training reduces the effective volume ($\partial_{t}\mathrm{Vol}(\mathcal{M})<0$), directly leading to the observed negative shift in semantic space.
@@ 3.1 The Limitations of Pixel Space
Euclidean distance in pixel space is sensitive to translation and color shifts, potentially leading to ”shortcuts” where the model generates mean-mode artifacts (shadows) to minimize distance.
@@ 3.2 Feature Space Percolation
To address this, we extend our framework to the Semantic Space using pretrained networks (e.g., VGG-16).
@@
where $\phi(\cdot)$ extracts deep features. Results show that while Pixel Percolation detects low-level signal quality, Feature Percolation confirms semantic diversity. A robust model must satisfy $\Delta\varepsilon_{c}\approx 0$ in both views.
@@ 4.1 Theoretical Justification of the Topological Loss
Let $d_{(k)}^{\text{fake}}$ and $d_{(k)}^{\text{real}}$ be the $k$-th smallest pairwise (distances in a minibatch of size $B$) computed in a fixed metric space (e.g. VGG feature space). Minimizing the squared $\ell_{2}$ loss on the ordered spectra
@@
For fixed sample size $N$ and intrinsic dimension $d$, Theorem 1 implies $\varepsilon_{c}\propto\bigl(\mathrm{Vol}(\mathcal{M})\bigr)^{1/d}$. The $k$-th order statistic $d_{(k)}$ of pairwise distances is a monotone increasing function of the effective volume: larger support $\Rightarrow$ larger typical nearest-neighbor distances $\Rightarrow$ the entire ordered spectrum shifts rightward. The quadratic penalty is strictly convex in each $d_{(k)}^{\text{fake}}$ and has positive gradient $\frac{\partial\mathcal{L}_{\text{topo}}}{\partial d_{(k)}^{\text{fake}}}\propto 2(d_{(k)}^{\text{fake}}-d_{(k)}^{\text{real}})$. In the under-covering regime (typical after a few epochs, $d_{(k)}^{\text{fake}}<d_{(k)}^{\text{real}}$ for most $k$), the gradient is uniformly positive, repelling samples and increasing local volumes, which by Theorem 1 directly raises $\varepsilon_{c}$. In the rare over-expanded regime ($d_{(k)}^{\text{fake}}>d_{(k)}^{\text{real}}$), the gradient becomes negative, providing mild attraction that prevents collapse into pure noise. Thus the loss acts as a bidirectional topological regularizer. ∎
@@
Empirically, we observe the baseline settles in the under-covering regime ($\Delta\varepsilon_{c}\approx-1.1$, Table 1). Consequently, $\mathcal{L}_{\text{topo}}$ exerts a consistent repulsive force throughout late training, explaining the observed transition from $\Delta\varepsilon_{c}<0$ to $\Delta\varepsilon_{c}>0$ (Hyper-Generalization) in Section 4.3.

### v2

# Manifold Percolation: from generative model to Reinforce learning Metrics, Dynamics, and Topological Supervision

@@ Abstract
Generative modeling is typically framed as learning mapping rules, but from an observer’s perspective without access to these rules, the task manifests as disentangling the geometric support from the probability distribution. We propose that Continuum Percolation is uniquely suited for this support analysis, as the sampling process effectively projects high-dimensional density estimation onto a geometric counting problem on the support. In this work, we establish a rigorous isomorphism between the topological phase transitions of Random Geometric Graphs and the underlying data manifold in high-dimensional space. By analyzing the relationship between our proposed Percolation Shift metric and FID, we demonstrate that our metric captures structural pathologies (such as implicit mode collapse) where statistical metrics fail. Finally, we translate this topological phenomenon into a differentiable loss function to guide training. Experimental results confirm that this approach not only prevents manifold shrinkage but fosters a "Synergistic Improvement," where topological stability acts as a prerequisite for sustained high fidelity in both static generation and sequential decision-making.
1 Introduction
1.1 Motivation: The "Phantom Manifold" Thought Experiment
Why do we need a topological "Observer’s View" for generative modeling? Consider the following thought experiment.
Let $\mathcal{D}_{\text{noise}}=\{\bm{z}^{(k)}\}_{k=1}^{N}$ be a dataset of $N$ samples drawn from pure white noise, $\bm{z}\sim\mathcal{N}(0,\mathbf{I}_{D})$. By definition, the pixels are independent: the ground-truth Jacobian is strictly diagonal, $J_{ij}=\frac{\partial z_{i}}{\partial z_{j}}=0$ for $i\neq j$. The true data manifold is the entire ambient space $\mathbb{R}^{D}$ (or a high-dimensional ball), with no low-dimensional structure.
Now, suppose we train a powerful neural network (e.g., a diffusion model[1, 2]) to model this dataset. Due to the finite $N$ and the over-parameterized nature of deep networks, the model will inevitably overfit. It will memorize the specific noise patterns in $\mathcal{D}_{\text{noise}}$. Mathematically, the learned function $f_{\theta}$ will encode spurious correlations, resulting in a non-zero sensitivity map:
$\left|\frac{\partial\hat{x}_{i}}{\partial\hat{x}_{j}}\right|>0\quad\text{for }i\neq j.$ (1)
To the model, the data appears to lie on a complex, low-dimensional "Phantom Manifold" defined by these spurious dependencies.
Here lies the paradox:
  • From the Model’s Internal View (Loss): The training loss approaches zero. The model believes it has perfectly learned the "structure" of the data.
  • From the Physical View (Independence): The model is hallucinating structure where there is only entropy.
This paradox reveals that we cannot trust the model’s internal proxies (loss, likelihood) to judge geometric validity. We require an external, observer-centric metric capable of distinguishing between a Phantom Manifold (memorized noise clumps) and a True Manifold (semantic structure). This motivates our use of Continuum Percolation: a topological tool specifically designed to analyze the connectivity and volume of point clouds without accessing the generating rules.
@@ 2 Theoretical Framework: Manifold Percolation
In this section, we establish the mathematical foundation of our framework. We bridge the gap between generative modeling and statistical physics by formalizing the "Observer’s View" through the lens of Continuum Percolation Theory.
@@ 2.3 Scaling Laws and Intrinsic Dimension
Solving for $\varepsilon_{c}$ at $\langle k\rangle=\lambda_{c}$ yields $\varepsilon_{c}\propto(\int p^{2}(x)dx)^{-1/d}N^{-1/d}$. This confirms that the scaling law is robust to density variations, with the "effective volume" determined by the Rényi entropy. ∎
@@ 2.5 Metric Spaces: From Pixel to Semantic Percolation
While the standard RGG is constructed in the raw pixel space $\mathcal{X}\subset\mathbb{R}^{D}$, relying solely on Euclidean distance ($d_{\text{pix}}(x,y)=\|x-y\|_{2}$) can be misleading. Pixel-wise distance is known to be sensitive to translation and noise, often failing to capture high-level structural differences.
To rigorously evaluate the semantic topology, we extend the percolation framework to the deep feature space defined by a pre-trained encoder $\phi:\mathcal{X}\to\mathbb{R}^{d^{\prime}}$. We employ a VGG-16 [11]network pre-trained on ImageNet as the feature extractor $\phi$. The Semantic Percolation Process is defined by the edge set:
$E_{\varepsilon}^{\text{feat}}=\{(x^{(i)},x^{(j)})\mid\|\phi(x^{(i)})-\phi(x^{(j)})\|_{2}\leq\varepsilon\}.$ (12)
This projection allows us to detect "Semantic Collapse," where a model might maintain pixel variance (noise) but fails to cover the semantic diversity of the true manifold.
@@ 3 Manifold Dynamics and Topological Diagnosis
$\Delta\varepsilon_{c}=\varepsilon_{c}(\text{Model})-\varepsilon_{c}(\text{Real}).$ (13)
According to the Manifold Shrinkage Theorem (Theorem 1), a negative shift ($\Delta\varepsilon_{c}<0$) serves as a monotone proxy for manifold shrinkage, i.e., it provides strong evidence that
$\mathrm{Vol}(\mathcal{M}_{\text{model}})<\mathrm{Vol}(\mathcal{M}_{\text{real}}).$
@@ 3.1 The Geometric Mechanism of Shrinkage
To validate our theoretical framework, we first conduct a controlled experiment on a synthetic 2D manifold. We simulate Implicit Mode Collapse by reducing the intra-mode variance of a Gaussian mixture while preserving the mode centers. As predicted, the percolation phase transition curve shifts to the left, resulting in a negative $\Delta\varepsilon_{c}$. This confirms that our metric is sensitive to the effective volume of the support, effectively distinguishing between a dispersed (healthy) manifold and a contracted (collapsed) one.
@@
Figure 2: Controlled Experiment on a Toy Manifold. (a) Simulating Implicit Mode Collapse by reducing variance (red) relative to ground truth (gray). (b) The percolation curve shifts left ($\Delta\varepsilon_{c}<0$), empirically validating the shrinkage behavior predicted by Theorem 1.
3.2 The Deceptive Nature of Visual Projections
A common practice in generative modeling is to rely on dimensionality reduction (e.g., UMAP) for visual inspection. However, by contrasting visual projections with quantitative percolation curves, we uncover a critical Visual Illusion.
As shown in Figure 3a, UMAP projections depict the generated samples converging towards the data manifold, creating an impression of structural stability from Epoch 60 to 300. The projection at Epoch 300 appears visually dense and well-structured, potentially leading observers to believe the model has converged to a healthy state.
However, this visualization is geometrically unfaithful. It fails to capture the drastic changes in effective support volume. The true topological dynamics are revealed in Figure 3b (Percolation Curves):
  • Initial State (Epoch 0): The percolation curve is positioned far to the right of the real data baseline, corresponding to a large positive shift ($\Delta\varepsilon_{c}=+11.49$). This quantifies the initial high-entropy noise state (Zone III in our phase diagram, i.e., an over-expanded manifold), whereas UMAP merely displays this as a diffuse cloud without indicating the scale of the volume excess.
  • Late-Stage Collapse (Epoch 300): While UMAP suggests a robust manifold similar to earlier epochs, Fig. 3b reveals that the percolation curve has shifted significantly to the left of the baseline ($\Delta\varepsilon_{c}=-1.12$). This indicates Implicit Mode Collapse (Zone I: a contracted, low-volume manifold).
This discrepancy demonstrates that dimensionality reduction algorithms, by optimizing for local neighborhood preservation, tend to normalize density differences. They artificially “inflate” the collapsed, low-volume manifold (Ep 300) to fill the visualization space, rendering the shrinkage invisible to the naked eye.
Remark 1 (The Ambiguity of Visual Clumping: Collapse vs. Divergence).
Our experiments further identify a paradoxical Visual Clump phenomenon. When a model generates high-frequency noise or artifacts (Manifold Divergence in Zone III), $\Delta\varepsilon_{c}$ correctly exhibits a large positive spike ($\Delta\varepsilon_{c}\gg 0$). However, UMAP often projects these high-dimensional isotropic noise samples as a tight, off-manifold clump, visually indistinguishable from a collapsed mode. Thus, visual inspection is doubly deceptive: it conflates Implicit Shrinkage (Zone I) and Manifold Explosion (Zone III) into similar visual blobs. In contrast, our Percolation Shift cleanly separates them at the sign level in practice ($\Delta\varepsilon_{c}<0$ vs. $\Delta\varepsilon_{c}>0$), providing a robust diagnostic signal that visual projections lack.
(a) The Illusion of Stability: UMAP projections (top row) show the samples converging to the manifold location but mask the volume shrinkage. Visually, Ep 300 (red) appears as robust as Ep 100, contradicting the topological reality.
(b) The Topological Truth: Pixel Space (left) overlaps perfectly (illusion). However, Semantic Space (right) reveals the true dynamics: from Over-expansion at Ep 0 (curve right) to Shrinkage at Ep 300 (curve left).
(c) Quantitative Diagnosis: The dual-axis plot confirms the collapse. While FID (red) improves, $\Delta\varepsilon_{c}$ (blue) plummets, illustrating a Fidelity–Topology trade-off.
Figure 3: The Anatomy of Implicit Mode Collapse. A unified diagnostic framework. We use the rigorous topological signals from (b) and (c) to expose the visual illusions present in (a).
3.3 The Illusion of Convergence: Metric Decoupling
Beyond global statistics, we isolate a critical phase transition in the Baseline model to demonstrate the decoupling between statistical fidelity (FID) and topological health ($\Delta\varepsilon_{c}$).
Case Study: The Trap of Epoch 270.
We compare two checkpoints of the Baseline model: Epoch 180 and Epoch 270. As shown in Table 1, a practitioner relying solely on FID would mistakenly prefer Epoch 270.
Table 1: Conflict between Fidelity and Topology. Comparing two training stages. While FID improves (lowers) at Epoch 270, the Percolation Shift reveals a hidden structural collapse.
@@ 3.3 The Illusion of Convergence: Metric Decoupling
Model State FID $\downarrow$ Shift ($\Delta\varepsilon_{c}$) $\uparrow$ Diagnosis
Epoch 180 25.74 -0.19 Topological Peak
Epoch 270 25.02 -1.15 Hidden Fracture
However, the Percolation Shift reveals a drastic degradation ($\Delta\varepsilon_{c}$ drops from -0.19 to -1.15). This implies that the model has begun to contract onto a subset of modes to optimize likelihood, sacrificing diversity for a marginal gain in precision.
3.4 Hierarchical Optimization Dynamics
Diagnostic Tool: Latent Interpolation.
Beyond global statistics, we examine how the learned manifold behaves along interpolation paths between samples. For each row, we linearly interpolate between two endpoints A and B in the latent space, a technique central to Latent Diffusion Models [3], and decode the intermediate points. A disconnected manifold will manifest as discontinuous jumps or artifacts along this path.
(a) Epoch 180 (Structural Peak shift=-0.19 FID=25.74): The interpolation path is smooth and continuous. The model has learned the connected topology of the data manifold.
(b) Epoch 270 (Hidden Fracture shift=-1.15 FID=25.02): Despite better FID, the interpolation path reveals "holes" and rapid transitions (e.g., undefined artifacts in middle frames). The manifold has fractured into disjoint high-density islands.
Figure 4: Visualizing the Topological Fracture. Comparing latent space interpolations at the two key epochs. The degradation in connectivity (b) is invisible to FID but perfectly captured by the negative Percolation Shift.
The visual evidence in Figure 4 confirms the mechanism of decoupling using this technique. At Epoch 270, the model has "over-polished" the high-density regions (lowering FID) but severed the bridges between them (lowering Shift).
This suggests that generative optimization operates on a hierarchy:
  • First-Order (FID/MSE): Measures point-wise quality. It improves by contracting variance.
  • Second-Order (Topology): Measures connectivity and shape. It degrades when contraction breaks the manifold structure.
Standard training greedily optimizes the first-order objective, often at the expense of second-order integrity. The Percolation Shift acts as a necessary "second-derivative test" to detect this overfitting.
3.5 Cross-Domain Dynamics: GANs vs. RL vs. Diffusion
The phenomenon of Manifold Shrinkage is not unique to diffusion models; we posit that it serves as a universal signature for generative pathology across different learning paradigms. To verify this, we extended our topological tracking to Generative Adversarial Networks (GANs) and Reinforcement Learning (RL).
Figure 5: The Topological Signature of GAN Training. We tracked the Percolation Shift $\Delta\varepsilon_{c}$ during standard GAN training. The curve rises from initialization but stabilizes in a persistent negative regime ($\Delta\varepsilon_{c}\approx-1.5$, shaded red region). Unlike diffusion models which may initially over-expand and then shrink, this confirms that GANs suffer from inherent Manifold Shrinkage, providing rigorous topological evidence for the classical “Mode Dropping” phenomenon.
As illustrated in Figure 5, distinct optimization objectives leave characteristic topological fingerprints:
  • GANs (Adversarial Shrinkage): GANs consistently oscillate in the negative regime ($\Delta\varepsilon_{c}<0$). The discriminator forces the generator to focus only on the highest-density modes to survive, leading to permanent “Mode Dropping” where the generated manifold volume remains strictly smaller than the true data support.
  • Standard RL (Policy Collapse): As we will demonstrate in Section 6, standard Policy Gradient methods (like PPO) exhibit even more extreme shrinkage. The shift $\Delta\varepsilon_{c}$ drifts deep into the negative regime ($\Delta\varepsilon_{c}\to$-$\infty$) as the policy converges onto a single near-deterministic trajectory, sacrificing exploration for exploitation.
  • Topo-Regularized Models: In contrast, our proposed topological supervision acts as a stabilizer, maintaining $\Delta\varepsilon_{c}\approx 0$. It effectively counteracts the shrinkage forces of adversarial or reward-based objectives, balancing fidelity with entropic diversity.
@@ 4.1 The Limitations of Pixel Space
Euclidean distance in pixel space is sensitive to translation and color shifts, potentially leading to "shortcuts" where the model generates mean-mode artifacts (shadows) to minimize distance.
@@ 4.2 Feature Space Percolation
To address this, we extend our framework to the Semantic Space using pretrained networks (e.g., VGG-16). The semantic distance is defined as:
@@
where $\phi(\cdot)$ extracts deep features.
Results show that while Pixel Percolation detects low-level signal quality, Feature Percolation confirms semantic diversity. As illustrated in Figure 6, standard models hit a "Topological Glass Ceiling," failing to fully cover the semantic manifold. A robust model must satisfy $\Delta\varepsilon_{c}\approx 0$ in both views.
Figure 6: Evolution of Semantic Percolation Shift ($\Delta\varepsilon_{c}$) over Training Epochs. Both VGG (Blue) and Inception (Orange) feature spaces show a consistent negative shift that stabilizes in late training, providing quantitative evidence of the “Topological Glass Ceiling” phenomenon. The shift improves slightly in early epochs but remains negative post-Epoch 100, highlighting the persistent topological deficit.
@@ 5.1 Theoretical Justification of the Topological Loss
Let $d_{(k)}^{\text{fake}}$ and $d_{(k)}^{\text{real}}$ be the $k$-th smallest pairwise distances in a minibatch of size $B$, computed in a fixed metric space (e.g. VGG feature space). Minimizing the squared $\ell_{2}$ loss on the ordered spectra
@@
For fixed sample size $N$ and intrinsic dimension $d$, Theorem 1 and the scaling law imply $\varepsilon_{c}\propto\bigl(\mathrm{Vol}(\mathcal{M})\bigr)^{1/d}$. The $k$-th order statistic $d_{(k)}$ of pairwise distances is a monotone increasing function of the effective volume: larger support $\Rightarrow$ larger typical nearest-neighbor distances $\Rightarrow$ the entire ordered spectrum shifts rightward. The quadratic penalty is strictly convex in each $d_{(k)}^{\text{fake}}$ and has positive gradient $\frac{\partial\mathcal{L}_{\text{topo}}}{\partial d_{(k)}^{\text{fake}}}\propto 2(d_{(k)}^{\text{fake}}-d_{(k)}^{\text{real}})$. In the under-covering regime (typical after a few epochs, $d_{(k)}^{\text{fake}}<d_{(k)}^{\text{real}}$ for most $k$), the gradient is uniformly positive, repelling samples and increasing local volumes, which by Theorem 1 directly raises $\varepsilon_{c}$. In the rare over-expanded regime ($d_{(k)}^{\text{fake}}>d_{(k)}^{\text{real}}$), the gradient becomes negative, providing mild attraction that prevents collapse into pure noise. Thus the loss acts as a bidirectional topological regularizer. ∎
@@
Empirically, we observe the baseline settles in the under-covering regime ($\Delta\varepsilon_{c}\approx-1.1$). Consequently, $\mathcal{L}_{\text{topo}}$ exerts a consistent repulsive force throughout late training, explaining the observed transition from $\Delta\varepsilon_{c}<0$ to $\Delta\varepsilon_{c}>0$ (Hyper-Generalization) in Section 4.3.